# Efficient Bayesian analysis of kilonovae and GRB afterglows with fiesta

### v1

@@ 4.2 Performance
Besides the cost of the likelihood function, another important factor that influences the time needed to sample the posterior is the dimension of the parameter space. The parameter space at least contains the base model parameters listed in Table 1, but can be extended with parameters to model the systematic uncertainty. As mentioned above, in fiesta, the systematic uncertainty $\sigma_{\text{sys}}$ can either be set to a constant value or it can be constructed from a set of sampling parameters. By introducing these nuisance parameters $\sigma^{(k)}_{\rm sys}(f)$, the systematic uncertainty can even become time- and filter-dependent through Eq. (3). However, when adding parameters to the sampler, this will impact the sampling time. In Fig. 7, we show how the sampling time (i.e. total runtime minus just-in-time compilation, which remains roughly constant over the cases considered here) per effective sample size of a fiesta inference evolves when more systematic nuisance parameters are added. While initially the sampling time per effective sample increases notably when going from 10 to 20 sampling dimensions, at higher dimensionality, the increase remains limited. This is opposite to the behavior of conventional samplers. For instance, in Table II of Jhawar et al. (2025) it is shown that when using the PyMultinest nested sampler in nmma, the sampling time increases from $11$ minutes when sampling $6$ parameters to $107$ minutes for a posterior of dimension $21$. We of course note that a setting with more than $20$ nuisance parameters seems superfluous for real-life data analysis, however, Fig. 7 emphasizes the capability of the flowMC sampler to handle large parameter spaces efficiently. This is also useful when combining different models at once, e.g., for joint analysis of GRB afterglow and KN emission.
@@ 5 Applications
In this section we apply our newly developed inference framework to two instances of real observations, namely AT2017gfo/GRB170817A, and GRB211112A. We use the best-performing surrogates from Sec. 3, i.e., the FluxModel cVAE’s for the GRB afterglow surrogates and the FluxModel with MLP architecture for the KN surrogate, to jointly analyze the KN emission and GRB afterglow emission in these events.
@@ 5.1 AT2017gfo/GRB170817A
Figure 8: Posterior of the joint KN+GRB afterglow analyses of AT2017gfo/GRB170817A. Selected parameters are shown in the corner plot. The full corner plot can be accessed in our data repository. The lightblue contours indicate the posterior when the GRB afterglow part is fit with afterglowpy, magenta is the posterior from pyblastafterglow. Both inferences use the KN surrogate from possis.
@@
As mentioned above in Sec. 1, the most prominent instance of a KN and GRB afterglow occurring together is the electromagnetic counterpart to the BNS merger associated with GW170817 (Abbott et al., 2017c, d). We use fiesta to reanalyze the joint lightcurve of these events, performing two analyses. One analysis uses our KN surrogate from possis together with the surrogate for the afterglowpy Gaussian jet, the other analysis uses the same KN surrogate but together with the pyblastafterglow Gaussian jet surrogate. Our priors are uniform in the ranges specified in Table 1, except for the inclination where we set $\iota\sim\mathcal{U}(0,\pi/4)$ to avoid a second mode at $\iota\approx\pi/2$. We confirmed that this second mode is not an artifact from the surrogate, but instead a real solution of afterglowpy to the GRB170817A data, though an inclination $\iota>\pi/4$ seems implausible given the observation of the GRB prompt emission and the radio interferometry measurement of the jet inclination angle (Mooley et al., 2018; Finstad et al., 2018; Chen et al., 2019; Troja et al., 2019a). The handling of the systematic uncertainty is split between those filters containing the KN observations and those filters containing the GRB afterglow. In particular, we find that the systematic uncertainty around the KN model needs to be represented by four parameters spaced linearly across the KN time interval, each of which is sampled from a uniform prior $\mathcal{U}(0.5,2)$. On the other hand, for the GRB afterglow observations only a single systematic uncertainty parameter with a uniform prior $\mathcal{U}(0.3,2)$ is needed. The available KN data reach from 0.3 to 24 days, the GRB afterglow data span from 9 to 742 days. We only fit KN data points up to 10 days, since we find that later data points are not well-represented by the possis model. The reason for this behavior is potentially rooted in the breakdown of the local thermodynamical equilibrium in the by-then low-density environment (Waxman et al., 2019; Pognan et al., 2022), which causes the possis predictions to become less applicable. We confirmed that adding the available KN data points at $t>10$ days does not significantly alter the posterior. We fix the luminosity distance at $d_{L}=43.6$ Mpc and redshift at $z=0.009727$ (Chornock et al., 2017).
@@
In Fig. 8 we show the posterior of our joint KN+GRB afterglow inferences. For the GRB afterglow parameters, we find good agreement between the afterglowpy and pyblastafterglow model, as well as good agreement compared to previous analyses (Ryan et al., 2020; Pang et al., 2023; Ghirlanda et al., 2019; Troja et al., 2019a). The estimated isotropic kinetic energy from afterglowpy at $\log_{10}(E_{0})=54.0^{+1.7}_{-2.5}$ is higher than for pyblastafterglow $\log_{10}(E_{0})=52.2^{+2.3}_{-1.8}$, as is the ambient density with $\log_{10}(n_{\rm ism})={-0.29}^{+1.64}_{-2.86}$ from afterglowpy and $\log_{10}(n_{\rm ism})={-2.25}^{+1.9}_{-2.33}$ from pyblastafterglow. We attribute this difference to the fact that generally pyblastafterglow lightcurves are brighter than afterglowpy lightcurves due to the different microphysics and radiation scheme. Both analyses find a jet opening angle of about $\theta_{c}=4-10\,^{\circ}$ and the inclination $\iota=23^{\circ}-44^{\circ}$, all values quoted at the 95% credibility level. The inclination is in line or slightly above the value inferred from the super-luminal motion of the jet and other analyses (Mooley et al., 2018; Troja et al., 2019a; Mooley et al., 2022; Ryan et al., 2024), though notably above the inclination inferred from the GW data alone assuming the known luminosity distance from the host galaxy (Abbott et al., 2017a).
@@
The KN parameters offer a more peculiar picture than the afterglow parameters. While the parameters for the ejecta masses broadly agree with estimates from previous works (Anand et al., 2023; Breschi et al., 2024; Pang et al., 2023; Sarin et al., 2024), we find that, in particular, the electron fractions rail against extreme values to accommodate the lightcurve data points after 10 days. For both analyses, we find $\bar{Y}_{e,\text{dyn}}=0.33^{+0.02}_{-0.05}$, and $Y_{e,\text{wind}}=0.24^{+0.02}_{-0.02}$ at 95% credibility. These values contradict the general standard picture in which the dynamical ejecta is neutron-rich, i.e. $Y_{e,\text{dyn}}\lesssim 0.25$, and the polar wind ejecta has a higher electron fraction (Metzger & Fernández, 2014; Metzger, 2020; Shibata & Hotokezaka, 2019; Kasen et al., 2017; Nedora et al., 2021). It should be noted, however, that our parameter $\bar{Y}_{e,\text{dyn}}$ refers to the mass-averaged electron fraction according to the distribution from Eq. (2) in Anand et al. (2023). Hence, the dynamical ejecta would still contain some portion with low electron fraction. Nevertheless, this value seems high compared to the (uniform) electron fraction in the wind component. The posterior distributions also indicate very low values for $v_{\text{ej, wind}}$ at the prior edge of 0.05 $c$, which is in line with previous works (Breschi et al., 2021, 2024; Anand et al., 2023). By comparing the possis training and test data to the AT2017gfo lightcurve, we confirm that the aforementioned issues are not linked to the performance of our ML surrogate, but instead point toward a systematic issue. Specifically, it seems hard to reconcile the slow descent of the $g$ band with the steep decline of the infra-red magnitudes in the 2mass filters. This can be seen in Fig. 9 where we show the best-fit lightcurves from our analyses for selected filters. Some tension between the bluer and redder components in AT2017gfo has been noted for instance in Breschi et al. (2021) and Hussenot-Desenonges et al. (2025) as well. As mentioned, this might be related to the breakdown of local thermo-dynamical equilibrium in the late-time ejecta (Waxman et al., 2019; Pognan et al., 2022), or other systematic reasons such as dominance of a few nucleonic decays in heating process (Kasen & Barnes, 2019), unexpected opacity evolution (Kasen & Barnes, 2019; Tanaka et al., 2020; Pognan et al., 2022; Gillanders et al., 2024), or ejecta geometry (Collins et al., 2024; King et al., 2025) and emphasizes the need for further research on the early and late KN emission mechanisms.
@@ 5.2 GRB211211A
Figure 10: Posterior of the joint KN+GRB afterglow analyses of GRB211211A. Selected parameters are shown in the corner plot. The full corner plot can be accessed in our data repository The lightblue contours indicate the posterior when the GRB afterglow part is fit with afterglowpy, magenta is the posterior from pyblastafterglow. Both inferences use the KN surrogate from possis. The lightgrey contours show the analysis from Kunert et al. (2024) as a reference.
@@
Figure 11: Best-fit lightcurves from the joint KN+GRB afterglow analyses of GRB211211A for selected photometric filters. Layout as in Fig. 9. Upper detection limits are shown as triangles.
@@
GRB211211A was a long GRB observed relatively close through the Swift observatory (Rastinejad et al., 2022). Subsequent analyses of the near-optical and X-ray emission indicated that a KN component was contributing to the emission (Rastinejad et al., 2022; Troja et al., 2022; Mei et al., 2022; Yang et al., 2022; Kunert et al., 2024). We reanalyze the lightcurve data compiled in Kunert et al. (2024) with fiesta, using the same surrogates as above for AT2017gfo/GRB170817A. Since the time range of the possis surrogate is restricted to the timespan of 0.2–26 days and the validity of the possis base model might be restricted to an even smaller time range, we exclude early data points before 0.2 days in the X-ray and uvot filters. We set the priors uniform in the ranges from Table 1 and fix the luminosity distance at $d_{L}=358$ Mpc and the redshift at $z=0.0763$, using Planck18 cosmology (Aghanim et al., 2020). Unlike in AT2017gfo/GRB170817A, the GRB afterglow and KN emission are not featured separately in the lightcurve data and are superimposed in the UV, optical, and infrared due to the GRB afterglow being seen on axis. Therefore, for these filters, we setup four systematic uncertainty parameters spaced linearly between 0.2–10 days, sampled from a uniform prior $\mathcal{U}(0.5,2)$. For the radio and X-ray band, we set four separate systematic uncertainty parameter that cover the time range from 0.2 to 10 days, sampled from a uniform prior $\mathcal{U}(0.3,2)$. This time range is sufficient, even though the X-ray data contain a late data point at 150 days. However, this data point is an upper detection limit and is fulfilled by our fit regardless of the systematic uncertainty. The posteriors are shown in Fig. 10, the best-fit lightcurves in Fig. 11. As in the case for GRB170817A, the inferred GRB afterglow parameters do not differ significantly between the analyses with afterglowpy and pyblastafterglow. The energy from afterglowpy is again slightly higher as the energy inferred from pyblastafterglow, $\log_{10}(E_{0})={51.7}^{+2.3}_{-1.4}$ compared to $\log_{10}(E_{0})={51.2}^{+1.2}_{-1.0}$, which is also the case for the interstellar densities from afterglowpy $\log_{10}(n_{\rm ism})={0.5}_{-3.4}^{+1.4}$ compared to pyblastafterglow at $\log_{10}(n_{\rm ism})={1.4}_{-3.2}^{+0.6}$. The microphysical parameters $p$, $\epsilon_{e}$, and $\epsilon_{B}$ are consistent between afterglowpy and pyblastafterglow.
@@
The KN parameters are relatively unconstrained, the posterior is impaired by the lower prior bound on the ejecta masses. The KN parameters for the dynamical component, i.e., its velocity and electron fraction, essentially recover the prior. The wind ejecta parameters are somewhat better constrained, though still relatively broad with $v_{\rm ej,wind}={0.12}_{-0.05}^{+0.03}$ and $Y_{e,\text{wind}}={0.27}_{-0.07}^{+0.09}$. Despite this fact, we find that the KN component is likely needed to explain the data of GRB211211A, as a simple GRB afterglow analysis results in poor fits, in line with the findings based on Bayesian model selection from Kunert et al. (2024).
@@
In Fig. 10 we also show the results from the analysis of Kunert et al. (2024). Though this analysis also uses the afterglowpy Gaussian jet model, unlike our analysis it does not discard early data points before 0.2 days and relies on a different kilonova model (Bulla, 2019; Dietrich et al., 2020; Anand et al., 2021; Almualla et al., 2021) with different priors for ejecta masses. Additionally, Kunert et al. (2024) also uses an inclination prior that is dependent on the sampled jet opening angle. Hence, the inferred parameters slightly differ, most notably for the ejecta masses and the interstellar medium density. However, the total ejecta masses $m_{\rm ej}=m_{\rm ej,dyn}+m_{\rm ej,wind}$ between the two different kilonova models is consistent, since our analyses both result in $m_{\rm ej}={0.028}_{-0.012}^{+0.045}$ while the analysis of Kunert et al. (2024) finds $m_{\rm ej}={0.021}_{-0.011}^{+0.043}$. This indicates that even if different assumptions in KN modeling lead to quantitatively different results for the specific model parameters, general features and trends may still be well constrained.

### v2

@@ 4.2 Performance
Besides the cost of the likelihood function, the size of the parameter space also influences the sampling time. The parameter space at least contains the base model parameters listed in Table 3, but can be extended with parameters to model the systematic uncertainty. As mentioned above, in fiesta, the systematic uncertainty, $\sigma_{\text{sys}}$, can either be set to a constant value or it can be constructed from a set of sampling parameters. By introducing these nuisance parameters $\sigma^{(k)}_{\rm sys}(f)$, the systematic uncertainty can even become time- and filter-dependent through Eq. (5). However, when adding parameters to the sampler, this will impact the sampling time. In Fig. 7, we show how the sampling time (i.e. total runtime minus just-in-time compilation, which remains roughly constant over the cases considered here) per effective sample size of a fiesta inference evolves when more systematic nuisance parameters are added. While initially the sampling time per effective sample increases notably when going from 10 to 20 sampling dimensions, at a higher dimensionality, the increase remains limited. This is opposite to the behavior of conventional samplers. For instance, in Table II of Jhawar:2024ezm it is shown that when using the PyMultinest nested sampler in nmma, the sampling time increases from $11$ minutes when sampling $6$ parameters to $107$ minutes for a posterior of dimension $21$. We of course note that a setting with more than $20$ nuisance parameters seems superfluous for real-life data analysis, however, Fig. 7 emphasizes the capability of the flowMC sampler to handle large parameter spaces efficiently. This is also useful when combining different models at once, for instance, in joint analyses of GRB afterglow and KN emission.
@@ 5 Applications
In this section, we apply our newly developed inference framework to two instances of real observations, namely, AT2017gfo/GRB170817A and GRB211112A. We use the best-performing surrogates from Sect. 3, namely, the FluxModel cVAE’s for the GRB afterglow surrogates and the FluxModel with MLP architecture for the KN surrogate, to jointly analyze the KN emission and GRB afterglow emission in these events.
@@ 5.1 AT2017gfo/GRB170817A
Figure 8: Posterior of the joint KN+GRB afterglow analyses of AT2017gfo/GRB170817A. Selected parameters are shown in the corner plot. The full corner plot can be accessed in our data repository. The lightblue contours indicate the posterior where the GRB afterglow part is fitted with afterglowpy, magenta is the posterior from pyblastafterglow. Both inferences use the KN surrogate from possis.
@@
As mentioned in Sect. 1, the most prominent instance of a KN and GRB afterglow occurring together is the electromagnetic counterpart to the BNS merger associated with GW170817 (LIGOScientific:2017vwq; LIGOScientific:2017ync). We used fiesta to reanalyze the joint light curve of these events, performing two analyses. One analysis uses our KN surrogate from possis together with the surrogate for the afterglowpy Gaussian jet, while the other analysis uses the same KN surrogate but together with the pyblastafterglow Gaussian jet surrogate. Our priors are uniform within the ranges specified in Table 3, except for the inclination, where we set $\iota\sim\mathcal{U}(0,\pi/4)$ to avoid a second mode at $\iota\approx\pi/2$. We confirmed that this second mode is not an artifact from the surrogate, but instead a solution of afterglowpy to the GRB170817A data, although an inclination of $\iota>\pi/4$ seems implausible given the observation of the GRB prompt emission and the radio interferometry measurement of the jet inclination angle (Mooley:2018qfh; Ghirlanda:2018uyx; Mooley:2022uqa). The handling of the systematic uncertainty is split between those filters containing the KN observations and those filters containing the GRB afterglow. In particular, we find that the systematic uncertainty around the KN model needs to be represented by four parameters spaced linearly across the KN time interval, each of which is sampled from a uniform prior $\mathcal{U}(0.5,2)$. On the other hand, for the GRB afterglow observations only a single systematic uncertainty parameter with a uniform prior $\mathcal{U}(0.3,2)$ is needed. The available KN data reach from 0.3 to 24 days, the GRB afterglow data span from 9 to 742 days. We only fit KN data points up to 10 days, since we find that later data points are not well-represented by the possis model. The reason for this behavior is potentially rooted in the breakdown of the local thermodynamical equilibrium in the by-then low-density environment (Waxman:2019png; Pognan:2022pix), which causes the possis predictions to become less applicable. We confirmed that adding the available KN data points at $t>10$ days does not significantly alter the posterior. We fixed the luminosity distance at $d_{L}=43.6$ Mpc and redshift at $z=0.009727$ (Chornock:2017sdf).
@@
In Fig. 8, we show the posterior of our joint KN+GRB afterglow inferences. For the GRB afterglow parameters, we find good agreement between the afterglowpy and pyblastafterglow models, as well as good agreement compared to previous analyses (Ryan:2019fhz; Pang:2022rzc; Ghirlanda:2018uyx; Troja:2018uns). The estimated isotropic kinetic energy from afterglowpy at $\log_{10}(E_{0})=54.0^{+1.7}_{-2.5}$ is higher than for pyblastafterglow $\log_{10}(E_{0})=52.2^{+2.3}_{-1.8}$, as is the ambient density with $\log_{10}(n_{\rm ism})={-0.29}^{+1.64}_{-2.86}$ from afterglowpy and $\log_{10}(n_{\rm ism})={-2.25}^{+1.9}_{-2.33}$ from pyblastafterglow. All values are quoted at the 95% level. We attribute this difference to the fact that generally pyblastafterglow light curves are brighter than afterglowpy light curves due to the different microphysics and radiation scheme. Both analyses find a jet opening angle of about $\theta_{c}=4-10\,^{\circ}$ and the inclination $\iota=23^{\circ}-44^{\circ}$. This value is partially consistent with analyses that include the displacement of the apparent superluminal centroid to the multiband light curve data (Mooley:2018qfh; Ghirlanda:2018uyx; Mooley:2022uqa), where a range of $\iota\approx 14^{\circ}-28^{\circ}$ is inferred. Instead, more recent analyses also find $\iota=17.2^{\circ}-21.2^{\circ}$ (Govreen-Segal:2023wvs) and $\iota=18^{\circ}-24^{\circ}$ (Ryan:2023pzk), which is lower than our estimate. Yet, our credible interval remains consistent with $\iota\approx 0^{\circ}-40^{\circ}$ inferred from the GW data alone (LIGOScientific:2017adf), using current estimates for the Hubble constant.
@@
The KN parameters offer a more peculiar picture than the afterglow parameters. While the parameters for the ejecta masses broadly agree with estimates from previous works (Anand:2023jbz; Breschi:2024qlc; Pang:2022rzc; Sarin:2023khf), we find that, in particular, the electron fractions converge to rather extreme values. For both analyses, we find $\bar{Y}_{e,\text{dyn}}=0.33^{+0.02}_{-0.05}$, and $Y_{e,\text{wind}}=0.24^{+0.02}_{-0.02}$ at 95% credibility. These values contradict the general standard picture in which the dynamical ejecta are neutron-rich, i.e. $Y_{e,\text{dyn}}\lesssim 0.25$, and the polar wind ejecta has a higher electron fraction (Metzger:2014ila; Metzger:2019zeh; Shibata:2019wef; Kasen:2017sxr; Nedora:2020hxc). It should be noted, however, that our parameter $\bar{Y}_{e,\text{dyn}}$ refers to the mass-averaged electron fraction according to the distribution from Eq. (2) in Anand:2023jbz. Hence, the dynamical ejecta would still contain some portion with low electron fraction. Nevertheless, this value seems high compared to the (uniform) electron fraction in the wind component. The posterior distributions also indicate very low values for $v_{\text{ej, wind}}$ at the prior edge of 0.05 $c$, which is in line with previous works (Breschi:2021tbm; Breschi:2024qlc; Anand:2023jbz). By comparing the possis training and test data to the AT2017gfo light curve, we confirm that the aforementioned issues are not linked to the performance of our ML surrogate, but instead point towards a systematic issue. Specifically, it seems hard to reconcile the slow descent of the $g$ band with the steep decline of the infrared magnitudes in the 2mass filters. This can be seen in Fig. 9 where we show the best-fit light curves from our analyses for selected filters. Some tension between the bluer and redder components in AT2017gfo has been noted for instance in Breschi:2021tbm and Hussenot-Desenonges:2025gik as well. As mentioned, this might be related to the breakdown of local thermodynamical equilibrium in the late-time ejecta (Waxman:2019png; Pognan:2022pix), or other systematic reasons such as dominance of a few nucleonic decays in the heating process (Kasen:2018drm), unexpected opacity evolution (Kasen:2018drm; Tanaka:2019iqp; Pognan:2022pix; Gillanders:2023jpd), or ejecta geometry (Collins:2023btn; King:2025tqo), and emphasizes the need for further research on the early and late KN emission mechanisms.
@@ 5.2 GRB211211A
Figure 10: Posterior of the joint KN+GRB afterglow analyses of GRB211211A. Selected parameters are shown in the corner plot. The full corner plot can be accessed in our data repository. The lightblue contours indicate the posterior when the GRB afterglow part is fit with afterglowpy, magenta is the posterior from pyblastafterglow. Both inferences use the KN surrogate from possis. The lightgrey contours show the analysis from Kunert:2023vqd as a reference.
@@
Figure 11: Best-fit light curves from the joint KN+GRB afterglow analyses of GRB211211A for selected photometric filters. Layout as in Fig. 9. Detection limits are shown as triangles.
@@
GRB211211A was a long GRB observed with the Swift observatory and later associated with a relatively nearby host at $\sim 350$ Mpc (Rastinejad:2022zbg). Subsequent analyses of the near-optical and X-ray emission indicated that a KN component was contributing to the emission (Rastinejad:2022zbg; Troja:2022yya; Mei:2022ncd; Yang:2022qmy; Kunert:2023vqd). We reanalyze the light curve data compiled in Kunert:2023vqd with fiesta, using the same surrogates as above for AT2017gfo/GRB170817A. Since the time range of the possis surrogate is restricted to the timespan of 0.2–26 days and the validity of the possis base model might be restricted to an even smaller time range, we exclude early data points before 0.2 days in the X-ray and uvot filters. We set the priors to be uniform within the ranges from Table 3 and fix the luminosity distance at $d_{L}=358$ Mpc and the redshift at $z=0.0763$, using Planck18 cosmology (Planck:2018vyg). Unlike in AT2017gfo/GRB170817A, the GRB afterglow and KN emission are not featured separately in the light curve data and are superimposed in the UV, optical, and infrared due to the GRB afterglow being seen on axis. Therefore, we sample four systematic nuisance parameters spaced linearly between 0.2–10 days, sampled from a uniform prior $\mathcal{U}(0.5,2)$. For the radio and X-ray band, we set four separate systematic uncertainty parameters that cover the time range from 0.2 to 10 days, sampled from a uniform prior $\mathcal{U}(0.3,2)$. This time range is sufficient, even though the X-ray data contain a late data point at 150 days. However, this data point is a detection limit and is fulfilled by our fit regardless of the systematic uncertainty. The posteriors are shown in Fig. 10, the best-fit light curves in Fig. 11. As in the case for GRB170817A, the inferred GRB afterglow parameters do not differ significantly between the analyses with afterglowpy and pyblastafterglow. The energy from afterglowpy is again slightly higher as the energy inferred from pyblastafterglow, $\log_{10}(E_{0})={51.7}^{+2.3}_{-1.4}$ compared to $\log_{10}(E_{0})={51.2}^{+1.2}_{-1.0}$, which is also the case for the interstellar densities from afterglowpy $\log_{10}(n_{\rm ism})={0.5}_{-3.4}^{+1.4}$ compared to pyblastafterglow at $\log_{10}(n_{\rm ism})={1.4}_{-3.2}^{+0.6}$. The microphysical parameters $p$, $\epsilon_{e}$, and $\epsilon_{B}$ are consistent between afterglowpy and pyblastafterglow.
@@
The KN parameters are relatively unconstrained and some of their marginalized posterior distributions reach down to the lower prior bound. The KN parameters for the dynamical component, i.e., its velocity and electron fraction, essentially recover the prior. The wind ejecta parameters are somewhat better constrained, though still relatively broad with $v_{\rm ej,wind}={0.12}_{-0.05}^{+0.03}$ and $Y_{e,\text{wind}}={0.27}_{-0.07}^{+0.09}$. Despite this fact, we find that the KN component is likely needed to explain the data of GRB211211A, as a simple GRB afterglow analysis results in poor fits. Specifically, the reduced chi-squared statistic $\chi^{2}/\nu$ is 0.24 (0.28) in the joint inference with afterglowpy (pyblastafterglow) and the KN model, which increases to 1.0 (0.98) when the KN contribution is not included. This is in line with the findings based on Bayesian model selection from Kunert:2023vqd.
@@
In Fig. 10 we also show the results from the analysis of Kunert:2023vqd. Though this analysis also uses the afterglowpy Gaussian jet model, unlike our analysis it does not discard early data points before 0.2 days and relies on a different kilonova model (Bulla:2019muo; Dietrich:2020efo; Anand:2020eyg; Almualla:2021znj) with different priors for ejecta masses. Additionally, Kunert:2023vqd also uses an inclination prior that is dependent on the sampled jet opening angle. Hence, the inferred parameters slightly differ, most notably for the ejecta masses and the interstellar medium density. However, the total ejecta masses $m_{\rm ej}=m_{\rm ej,dyn}+m_{\rm ej,wind}$ obtained from the two different kilonova models are consistent. Our analyses both result in $m_{\rm ej}={0.028}_{-0.012}^{+0.045}$, while the analysis of Kunert:2023vqd finds $m_{\rm ej}={0.021}_{-0.011}^{+0.043}$. This indicates that even if different assumptions in KN modeling lead to quantitatively different results for the specific model parameters, the total ejecta mass can be inferred consistently.